# Avoid Wasted Annotation Costs in Open-set Active Learning with Pre-trained Vision-Language Model

### v1

@@ Abstract
Active learning (AL) aims to enhance model performance by selectively collecting highly informative data, thereby minimizing annotation costs. However, in practical scenarios, unlabeled data may contain out-of-distribution (OOD) samples, leading to wasted annotation costs if data is incorrectly selected. Recent research has explored methods to apply AL to open-set data, but these methods often require or incur unavoidable cost losses to minimize them. To address these challenges, we propose a novel selection strategy, CLIPN for AL (CLIPNAL), which minimizes cost losses without requiring OOD samples. CLIPNAL sequentially evaluates the purity and informativeness of data. First, it utilizes a pre-trained vision-language model to detect and exclude OOD data by leveraging linguistic and visual information of in-distribution (ID) data without additional training. Second, it selects highly informative data from the remaining ID data, and then the selected samples are annotated by human experts. Experimental results on datasets with various open-set conditions demonstrate that CLIPNAL achieves the lowest cost loss and highest performance across all scenarios. Code is available at https://github.com/DSBA-Lab/OpenAL.
@@ 1 Introduction
Deep learning has succeeded in various tasks such as classification and regression[7]. However, this success relies heavily on well-refined large datasets and the associated annotation costs. In practice, spending significant resources on annotating a large volume of data is difficult. To address this issue, research on active learning (AL) has been continuously studied, which aims to enhance model performance by selectively annotating informative data from unlabeled data[23].
@@
AL uses a query strategy to decide whether to annotate an unlabeled sample for model training. Traditional AL research, termed standard AL[18, 21, 15, 20], assumes that the unlabeled data is collected from the same domain as the training data called in-distribution data (ID). As shown in Fig. 2 (a), the query strategy of standard AL selects an unlabeled sample based on its informativeness, which refers to the degree to which the sample can help improve model performance. However, it is difficult to assume that unlabeled data only contains ID samples. For instance, if the classification task involves distinguishing between dogs and cats, it is assumed that the unlabeled data only contains images of dogs and cats. However, diverse data collected online might include other animals like guinea pigs or zebras (Fig. 1). If the unlabeled data includes OOD samples and only informativeness is considered, the selected data might include OOD samples, which cannot be used for training, leading to wasted annotation costs
Recent research has focused on open-set AL, which assumes that unlabeled data might contain out-of-distribution (OOD) data as well [4, 11, 10, 13]. Open-set AL considers not only the informativeness but also the purity, indicating whether the data is relevant to the domain of the training data. As illustrated in Fig. 2 (b), data selection should be based on both purity and informativeness to avoid selecting irrelevant OOD data in open-set AL assumption.
Figure 2: Comparison of data selection criterion for different active learning scenarios: (a) uses informativeness as the selection criteria for standard AL, while (b) and Ours use both informativeness and purity as the selection criteria for open-set AL.
Recent methods proposed for open-set AL consider both purity and informativeness. These methods can be categorized into those that perform contrastive learning on unlabeled data[4, 11] and those that train an OOD detector using collected OOD data[10, 13]. However, both approaches face cost inefficiency issues. Contrastive learning methods often fail to satisfy both purity and informativeness in all cases, leading to annotation cost losses. OOD detector methods require collecting OOD data to ensure high performance, inherently causing cost losses. Thus, while these methods perform better than standard AL, they still result in significant annotation cost losses. Fig. 3 shows the annotation costs wasted when using different active learning strategies.
Recent studies have explored using CLIP for zero-shot OOD detection[5, 9, 19]. To address these cost-inefficiency issues, we suggest a two-step query strategy using CLIP named CLIPN for Active Learning (CLIPNAL) for open-set data, as shown in Fig. 2 (Ours). In the first step, we use a pre-trained CLIP[12] to perform zero-shot OOD detection to evaluate unlabeled data’s purity. We utilize a pre-trained CLIPN[19] with an additional text encoder to incorporate the concept of “no” to evaluate the purity of unlabeled data. However, as textual information alone cannot fully represent all image classes, we enhance purity performance using visual similarity weights based on labeled ID data. Secondly, informativeness is assessed using the query strategy proposed in the Standard AL, which is only for the predicted ID data. Detecting OOD data with high accuracy makes it possible to convert open-set data into closed-set data. In other words, it leads to the same assumptions as Standard AL and helps solve the trade-off between purity and informativeness, a current issue in open-set AL.
The advantages of the proposed methodology can be summarized as follows: (1) It prevents annotation cost losses through high-performance OOD detection using pre-trained CLIPN and visual similarity weights reflecting ID data. (2) It does not require additional training for OOD detection because it uses a pre-trained model. (3) The OOD detection performance is not dependent on training, thus maintaining model robustness regardless of the proportion of OOD data in the unlabeled pool. Experiments assuming various open-set conditions demonstrate that our method outperforms both standard and existing open-set AL methods.
@@
Figure 3: Comparison of wasted annotation costs in relation to total annotation costs incurred when applying various AL methods to CIFAR10 and CIFAR100 configured as open-set data. $C_{\text{OOD}}$ represents the cost incurred when selected data is OOD data, and $C_{\text{ID}}$ represents the cost for ID data. Standard AL methods such as CONF[18], LL[21], and CORESET[15] result in higher cost losses compared to RANDOM due to the higher selection of OOD data. On the other hand, existing open-set AL methods like MQNet[11], EOAL[13], and LfOSA[10] show less cost loss than RANDOM but still suffer significant losses relative to the total annotation costs.
@@ 2.1 Standard Active Learning
The criteria for determining informativeness in query strategies can be broadly divided into two classes. The first is uncertainty-based query strategy[18, 21]. In a classification task, the uncertainty-based strategy assesses the degree of confidence in classifying unlabeled data based on the decision boundary of the trained model. If the model’s certainty about the data is low, it implies that the model finds it difficult to make an accurate judgment. Therefore, data with high uncertainty is selected for model training. The second is diversity-based query strategy[15, 20]. Despite the large data size, much of it contains redundant information. Hence, a small subset of data selected based on diversity can provide information similar to that of a larger dataset. Therefore, strategies that consider diversity select data that differ from the current data based on a comparison with labeled data.
@@ 2.2 Open-set Active Learning
Unlike the assumptions in standard AL, open-set AL considers the presence of OOD data in the unlabeled pool. Various methods have been proposed for this purpose. Firstly, CCAL[4] and MQNet[11] perform contrastive learning on unlabeled data to determine data purity using the trained model. CCAL uses SimCLR[1] and CSI[17] to reflect purity and informativeness, respectively, and proposes a joint query strategy that combines these scores for selecting data. However, contrastive learning alone focuses on purity, making it difficult to select highly informative data. To resolve this trade-off, MQNet adjusts the purity obtained from CSI and the informativeness from standard AL using meta-learning, allowing data collection to prioritize purity or informativeness as needed.
Figure 4: Differences in OOD detection between existing open-set active learning methods and CLIPNAL: Previous open-set active learning methods require datasets for training an OOD detector. LfOSA[10] and EOAL[13] even require the collection of OOD samples for training an OOD detector. However, CLIPNAL uses a pre-trained VLM that does not require datasets. Therefore, CLIPNAL reduces training costs and saves wasted annotation costs.
@@
Secondly, research has proposed training OOD detection models to assess data purity. LfOSA[10] trains a classifier to distinguish $K+1$ classes, including $K$ ID classes and one additional OOD class, using labeled ID and OOD data. The classifier predicts the class for each sample of unlabeled data, and based on the maximum activation value of the predicted class, a Gaussian Mixture Model (GMM) is trained with $K$ components. Each GMM assumes two Gaussian distributions. The distribution with the higher mean is assumed to represent the ID data distribution. Therefore, the higher the probability of a sample belonging to the Gaussian distribution with the higher mean, the higher its purity. Conversely, a lower probability indicates that the sample does not belong to any ID classes, meaning it is OOD class. However, LfOSA does not consider informativeness, limiting model performance. EOAL[13] addresses this issue by proposing a method to select data using a combined score based on two entropy measures representing purity and informativeness. Firstly, to assess informativeness, $K$ binary classifiers are trained, and the entropy scores for the unlabeled data are calculated using these $K$ binary classifiers corresponding to the ID classes. Secondly, to assess purity, a classifier is trained to classify $K+1$ classes, and the representations from this trained classifier for the labeled OOD data are used to perform clustering. The probability of belonging to each cluster is calculated based on the distances between clusters, and the entropy for data purity is computed for the unlabeled data. Combining these two scores, EOAL selects data with high purity and informativeness.
@@
While previous open-set AL methods show improved performance over standard AL by considering OOD data, they still incur annotation cost losses. MQNet, for instance, tends to select a large amount of OOD data because it prioritizes informativeness over purity when selecting high-informativeness data. Methods like LfOSA and EOAL require additional OOD data for OOD detector training, inherently leading to cost losses. Furthermore, both approaches necessitate extra training costs for collecting OOD data. However, our proposed method sequentially assesses purity and informativeness, avoiding the trade-off between these criteria. We use the pre-trained CLIP model for OOD detection, requiring no additional training (Fig. 4).
@@ 3 Proposed Method
Figure 5: Overview of evaluating the purity of an image in CLIPNAL. We utilize the CLIPN[19] to perform zero-shot OOD detection. According to CLIPN, two text encoders and an image encoder compute an image’s in-distribution (ID) probability. We additionally use visual similarity weights to update ID probability using labeled ID samples $S^{\text{ID}}$. Self-temperature tuning is used to find the optimal temperature parameter for improving zero-shot OOD performance.
The proposed CLIPNAL comprises two sequential processes: (1) measuring data purity based on a pre-trained CLIP and (2) computing informativeness of query strategies used in standard AL. In this section, we describe how our CLIP-based method assesses the purity of unlabeled data. As illustrated in Fig. 5, we use ID class names as linguistic information to calculate the ID probability of unlabeled images. Furthermore, we integrate visual information from labeled ID data to apply weighting to the ID probability. CLIPNAL process and an example of assessing purity can be found in Appendix. A.
@@ 3.1 Problem Statement
The open-set AL problem defined in this study aims to classify images into $K$ classes, with the goal of improving the performance of an image classifier $f^{\text{clf}}$ based on labeled data. We define the unlabeled data targeted for collection as open-set data $D_{\text{U}}=\{x_{i}\in\textbf{x}^{\text{ID}}\cup\textbf{x}^{\text{OOD}}\}_{i=1}^{N_{\text{U}}}$, assuming the presence of both ID and OOD data. In AL, a fixed budget $C$ is used for annotation costs in each round of data collection, repeated $R$ times. If annotating a single image costs one unit, setting $C$ to 10 means 10 images can be annotated per round. The selected queries from $D_{\text{U}}$ are denoted as $Q$. These queries are annotated by human experts, adding ID data $\textbf{x}^{\text{ID}}$ to $S^{\text{ID}}$ and OOD data to $S^{\text{OOD}}$. The labeled ID data $S^{\text{ID}}$ is used to train the classifier $f^{\text{clf}}$. Thus, open-set AL aims to maximize model performance by minimizing the proportion of OOD data and selecting highly informative data with minimal cost.
3.2 CLIPN
@@
Recently, large-scale datasets like MS COCO[8] and LAION-2B[14] have enabled CLIP[12] to be used not only for ID classification but also for OOD detection[5, 9, 19]. CLIPN[19], the latest method, enhances CLIP by adding a text encoder for the concept of “no”. Unlike previous CLIP-based zero-shot OOD detection methods[5, 9], which only consider proximity to class-specific text, CLIPN incorporates the notion of dissimilarity to achieve high performance in OOD detection. Therefore, our proposed method uses CLIPN to detect OOD data in the unlabeled pool.
@@
CLIPN assesses whether an image is OOD by considering the similarity between the image and class-specific textual information and the probability that the image does not belong to the class. First, the visual representation of an image $x_{i}^{\text{img}}$ is obtained by inputting it into the image encoder. Next, pre-defined textual templates with class names are input into the text encoder for each target class to calculate the average textual representation. Similarly, the “no” text encoder in CLIPN calculates the average representation for each class.
$z_{i}^{\text{img}}$ $=f^{\text{img}}(x_{i}^{\text{img}})$ (1)
$z_{j}^{\text{\text{text}}}$ $=\frac{1}{T}\sum_{k}^{T}{f^{\text{\text{text}}}(t^{\text{prompt}}_{k}(\text{class name}_{j}))}$
$z_{j}^{\text{\text{no text}}}$ $=\frac{1}{T}\sum_{k}^{T}{f^{\text{no text}}(t^{\text{prompt}}_{k}(\text{class name}_{j}))}.$
@@
Next, the probability that $x_{i}^{\text{img}}$ belongs to each class is calculated by computing the cosine similarity between the visual and textual representations and applying softmax.
$p_{ij}^{\text{yes}}=\frac{\exp{(z_{i}^{\text{img}}\cdot z_{j}^{\text{text}}/\tau)}}{\sum_{k}^{K}\exp{(z_{i}^{\text{img}}\cdot z_{k}^{\text{text}}/\tau)}},$ (2)
@@
where $\tau$ is a temperature parameter that adjusts model confidence. The probability that $x_{i}^{\text{img}}$ does not belong to each class $p_{ij}^{\text{no}}$ is defined as follows.
$p_{ij}^{\text{no}}=\frac{\exp{(z_{i}^{\text{img}}\cdot z_{j}^{\text{no text}}/\tau)}}{\exp{(z_{i}^{\text{img}}\cdot z_{j}^{\text{text}}/\tau)}+\exp{(z_{i}^{\text{img}}\cdot z_{j}^{\text{no text}}/\tau)}}.$ (3)
@@
$p_{ij}^{\text{ID}}$ $=p_{ij}^{\text{yes}}(1-p_{ij}^{\text{no}})$ (4)
$p_{i}^{\text{OOD}}$ $=1-\sum_{j}^{K}{p_{ij}^{\text{yes}}(1-p_{ij}^{\text{no}})}$

### v2

@@ Abstract
Active learning (AL) aims to enhance model performance by selectively collecting highly informative data, thereby minimizing annotation costs. However, in practical scenarios, unlabeled data may contain out-of-distribution (OOD) samples, which are not used for training, leading to wasted annotation costs if data is incorrectly selected. Therefore, to make active learning feasible in real-world applications, it is crucial to consider not only the informativeness of unlabeled samples but also their purity to determine whether they belong to the in-distribution (ID). Recent studies have applied AL under these assumptions, but challenges remain due to the trade-off between informativeness and purity, as well as the heavy dependence on OOD samples. These issues lead to the collection of OOD samples, resulting in a significant waste of annotation costs. To address these challenges, we propose a novel query strategy, VLPure-AL, which minimizes cost losses while reducing dependence on OOD samples. VLPure-AL sequentially evaluates the purity and informativeness of data. First, it utilizes a pre-trained vision-language model to detect and exclude OOD data with high accuracy by leveraging linguistic and visual information of ID data. Second, it selects highly informative data from the remaining ID data, and then the selected samples are annotated by human experts. Experimental results on datasets with various open-set conditions demonstrate that VLPure-AL achieves the lowest cost loss and highest performance across all scenarios.
keywords:
Active learning, Open-set active learning, Out-of-distribution detection, Vision-language models.
journal: Engineering Applications of Artificial Intelligence
[…]
[label1]organization=Department of Industrial Engineering, Seoul National University, country=Republic of Korea
@@ 1 Introduction
Active Learning (AL) is a research area proposed to improve model performance while minimizing the cost of data acquisition required for model training (Zhan et al., 2022). Although deep learning models have achieved impressive results on various tasks such as classification and regression, their success often depends on significant annotation costs (LeCun et al., 2015; Deng et al., 2009; Schuhmann et al., 2022). To address this issue, AL introduces diverse query strategies, designed to identify the most informative data samples from large pools of unlabeled data, thereby enhancing performance with minimal user cost (Wang and Shang, 2014; Yoo and Kweon, 2019; Sener and Savarese, 2018; Huang et al., 2022). AL has been extensively investigated and applied across a variety of fields such as action recognition (Li et al., 2023) and visual tracking (Yuan et al., 2023).
Traditional AL research, termed standard AL (Wang and Shang, 2014; Yoo and Kweon, 2019; Sener and Savarese, 2018; Yehuda et al., 2022), assumes that the unlabeled data is collected from the same domain as the training data called in-distribution data (ID). As shown in Figure 1 (a), the query strategy of standard AL selects an unlabeled sample focused on its informativeness, which refers to the degree to which the sample can help improve model performance by providing information that the model has not yet learned. However, it is difficult to assume that unlabeled data only contains ID samples. For instance, if the classification task involves distinguishing between dogs and cats, it is assumed that the unlabeled data only contains images of dogs and cats. However, diverse data collected online might include other animals like guinea pigs or zebras (Figure 2). If the unlabeled data includes out-of-distribution (OOD) samples and only informativeness is considered, the selected data might include OOD samples, which cannot be used for training, leading to wasted annotation costs. Therefore, to practically implement active learning, it is essential to consider the purity of the selected data from the unlabeled pool, ensuring that it shares the same distribution as the training data. In other words, it is important to minimize wasted annotation costs to ensure the collection of more in-distribution samples, which is essential for improving model performance.
Figure 1: Comparison of data selection criterion for different active learning scenarios: (a) uses informativeness as the selection criteria for standard AL, while (b) and Ours use both informativeness and purity as the selection criteria for open-set AL. Unlike (b), Ours evaluates purity and informativeness sequentially.
Recently, AL research has expanded to focus on the purity of unlabeled samples under the assumption that the unlabeled data pool contains not only ID samples but also OOD samples, as illustrated in Figure 1 (b), and it is called open-set AL (Du et al., 2021; Park et al., 2022; Ning et al., 2022; Safaei et al., 2024). Previous studies on open-set AL proposed query strategies that balance purity and informativeness through contrastive learning and meta-learning, adjusting the importance of these two metrics depending on the data collection stage to address their inherent dilemma (Du et al., 2021; Park et al., 2022). However, these methods face limitations, as they cannot simultaneously satisfy both purity and informativeness at every stage, resulting in an increasing proportion of OOD samples in the collected data. Another approach focuses on estimating the purity of unlabeled samples by training an OOD detector while simultaneously incorporating informativeness into the query strategy (Ning et al., 2022; Safaei et al., 2024). This approach also has limitations, as it requires a significant number of OOD samples to improve the performance of the OOD detector. Both approaches share a common drawback in that they require or heavily depend on collecting unnecessary data to satisfy the selection criteria for training samples.
We propose a sequential approach framework to address the aforementioned issues by assessing purity and then applying informativeness as shown in Figure 1. This sequential approach alleviates the trade-off problem and allows the flexible utilization of various query strategies proposed in standard AL while considering purity. However, achieving high purity accuracy is essential to minimize cost loss in the sequential process. To achieve high purity accuracy, we propose VLPure-AL, which leverages Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021) while reducing dependence on OOD samples.
@@
CLIP is trained to align representations of images and text in the same embedding space, enabling the classification of images using textual information. Due to these advantages, recent research has actively employed CLIP in the OOD detection domain (Esmaeilpour et al., 2022; Ming et al., 2022; Wang et al., 2023). In VLPure-AL, we enhance the process of assessing the purity of unlabeled samples by utilizing the textual information of CLIP’s text encoder and applying visual similarity weights, which leverage class-specific visual information from the collected labeled data. Furthermore, for more accurate purity assessment, we refine the method by incorporating unavoidable OOD samples through self-temperature tuning, which optimizes the temperature scaling factor used in the purity assessment process.
The advantages of our proposed method can be summarized as follows:
  • 1.
    By sequentially evaluating purity with high accuracy and considering informativeness, it resolves the trade-off issue between purity and informativeness while allowing the flexible application of Standard AL methods.
  • 2.
    VLPure-AL reduces dependence on OOD samples by leveraging a pre-trained CLIP model and minimizes cost loss through accurate OOD detection using textual information, visual similarity weights, and self-temperature tuning.
  • 3.
    The proposed method demonstrates lower annotation cost loss and superior performance compared to both standard AL and open-set AL methods under various open-set conditions.
This paper is organized as follows: Section 2 provides a review of standard AL and open-set AL. Section 3 demonstrates the workflow of the proposed framework and introduces our proposed query strategy for open-set AL. Section 4 reports experimental results under various open-set conditions. Finally, Section 5 summarizes our contributions and discusses the limitations and future direction of our work.
@@
Figure 3: Comparison of wasted annotation costs in relation to total annotation costs incurred when applying various AL methods to CIFAR10 and CIFAR100 configured as open-set data. $C_{\text{OOD}}$ represents the cost incurred when selected data is OOD data, and $C_{\text{ID}}$ represents the cost for ID data. Standard AL methods such as CONF, LL, and CORESET result in higher cost losses compared to RANDOM due to the higher selection of OOD data. On the other hand, existing open-set AL methods like MQNet, EOAL, and LfOSA show less cost loss than RANDOM but still suffer significant losses relative to the total annotation costs.
@@ 2.1 Standard Active Learning
The criteria for determining informativeness in query strategies can be broadly divided into two classes. The first is uncertainty-based query strategy (Wang and Shang, 2014; Yoo and Kweon, 2019). In a classification task, the uncertainty-based strategy assesses the degree of confidence in classifying unlabeled data based on the decision boundary of the trained model. If the model’s certainty about the data is low, it implies that the model finds it difficult to make an accurate judgment. Therefore, data with high uncertainty is selected for model training. The second is diversity-based query strategy (Sener and Savarese, 2018; Yehuda et al., 2022). Despite the large data size, much of it contains redundant information. Hence, a small subset of data selected based on diversity can provide information similar to that of a larger dataset. Therefore, strategies that consider diversity select data that differ from the current data based on a comparison with labeled data.
@@ 2.2 Open-set Active Learning
Unlike the assumptions in standard AL, open-set AL considers the presence of OOD data in the unlabeled pool. Various methods have been proposed for this purpose. Firstly, CCAL (Du et al., 2021) and MQNet (Park et al., 2022) perform contrastive learning on unlabeled data to determine data purity using the trained model. CCAL uses SimCLR (Chen et al., 2020) and CSI (Tack et al., 2020) to reflect purity and informativeness, respectively, and proposes a joint query strategy that combines these scores for selecting data. However, contrastive learning alone focuses on purity, making it difficult to select highly informative data. To resolve this trade-off, MQNet adjusts the purity obtained from CSI and the informativeness from standard AL using meta-learning, allowing data collection to prioritize purity or informativeness as needed.
@@
Secondly, research has proposed training OOD detection models to assess data purity. LfOSA (Ning et al., 2022) trains a classifier to distinguish $K+1$ classes, including $K$ ID classes and one additional OOD class, using labeled ID and OOD data. The classifier predicts the class for each sample of unlabeled data, and based on the maximum activation value of the predicted class, a Gaussian Mixture Model (GMM) is trained with $K$ components. Each GMM assumes two Gaussian distributions. The distribution with the higher mean is assumed to represent the ID data distribution. Therefore, the higher the probability of a sample belonging to the Gaussian distribution with the higher mean, the higher its purity. Conversely, a lower probability indicates that the sample does not belong to any ID classes, meaning it is OOD class. However, LfOSA does not consider informativeness, limiting model performance. EOAL (Safaei et al., 2024) addresses this issue by proposing a method to select data using a combined score based on two entropy measures representing purity and informativeness. Firstly, to assess informativeness, $K$ binary classifiers are trained, and the entropy scores for the unlabeled data are calculated using these $K$ binary classifiers corresponding to the ID classes. Secondly, to assess purity, a classifier is trained to classify $K+1$ classes, and the representations from this trained classifier for the labeled OOD data are used to perform clustering. The probability of belonging to each cluster is calculated based on the distances between clusters, and the entropy for data purity is computed for the unlabeled data. Combining these two scores, EOAL selects data with high purity and informativeness.
@@
While previous open-set AL methods show improved performance over standard AL by considering OOD data, they still incur annotation cost losses. MQNet, for instance, tends to select a large amount of OOD data because it prioritizes informativeness over purity when selecting high-informativeness data. Methods like LfOSA and EOAL require additional OOD data for OOD detector training, inherently leading to cost losses. By leveraging a pre-trained CLIP, we reduce dependence on OOD samples, accurately assess purity, and consider informativeness, effectively addressing the limitations of previous studies.
@@ 3 Proposed Method
Figure 4: Overview of evaluating the purity of an image in VLPure-AL. We utilize the CLIPN to perform zero-shot OOD detection. According to CLIPN, two text encoders and an image encoder compute an image’s in-distribution (ID) probability. We additionally use visual similarity weights to update ID probability using labeled ID samples $S^{\text{ID}}$. Self-temperature tuning is used to find the optimal temperature parameter for improving zero-shot OOD performance.
Figure 5: Example of the data purity assessment process in VLPure-AL.
The proposed VLPure-AL comprises two sequential processes: (1) measuring data purity based on a pre-trained CLIP and (2) computing informativeness of query strategies used in standard AL. In this section, we describe how our CLIP-based method assesses the purity of unlabeled data. As illustrated in Figure 4, we use ID class names as linguistic information to calculate the ID probability of unlabeled images. Furthermore, we integrate visual information from labeled ID data to apply weighting to the ID probability. An example of VLPure can be seen in Figure 5.
@@ 3.1 Problem Statement
The open-set AL problem defined in this study aims to classify images into $K$ classes, to improve the performance of an image classifier $f^{\text{clf}}$ based on labeled data. We define the unlabeled data targeted for collection as open-set data $D_{\text{U}}=\{x_{i}\in\textbf{x}^{\text{ID}}\cup\textbf{x}^{\text{OOD}}\}_{i=1}^{N_{\text{U}}}$, assuming the presence of both ID and OOD data. In AL, a fixed budget $C$ is used for annotation costs in each round of data collection, repeated $R$ times. If annotating a single image costs one unit, setting $C$ to 10 means 10 images can be annotated per round. The selected queries from $D_{\text{U}}$ are denoted as $Q$. These queries are annotated by human experts, adding ID data $\textbf{x}^{\text{ID}}$ to $S^{\text{ID}}$ and OOD data to $S^{\text{OOD}}$. The labeled ID data $S^{\text{ID}}$ is used to train the classifier $f^{\text{clf}}$. Thus, open-set AL aims to maximize model performance by minimizing the proportion of OOD data and selecting highly informative data with minimal cost.
3.2 CLIP-based OOD Detection
@@
Recently, large-scale datasets like MS COCO (Lin et al., 2014) and LAION-2B (Schuhmann et al., 2022) have enabled CLIP (Radford et al., 2021) to be used not only for ID classification but also for OOD detection (Esmaeilpour et al., 2022; Ming et al., 2022; Wang et al., 2023). CLIPN (Wang et al., 2023), the latest method, enhances CLIP by adding a text encoder for the concept of “no”. Unlike previous CLIP-based zero-shot OOD detection methods (Esmaeilpour et al., 2022; Ming et al., 2022), which only consider proximity to class-specific text, CLIPN incorporates the notion of dissimilarity to achieve high performance in OOD detection. Therefore, our proposed method uses CLIPN to detect OOD data in the unlabeled pool.
@@
CLIPN assesses whether an image is OOD by considering the similarity between the image and class-specific textual information and the probability that the image does not belong to the class. First, the visual representation $f_{i}$ of an image $x_{i}$ is obtained by inputting it into the image encoder $\phi$. Next, pre-defined textual templates $t_{k,j}$ such as ”A photo of $\{\text{class}_{j}\}$” are input into the text encoder $\psi$ for each target class to calculate the average textual representation $g_{j}$. Similarly, the “no” text encoder in CLIPN calculates the average representation $g_{j}^{no}$ for each class.
$f_{i}$ $=\phi^{\text{img}}(x_{i})$ (1)
$g_{j}$ $=\frac{1}{T}\sum_{l=1}^{T}\psi(t_{l,j})$
$g_{j}^{\text{\text{no}}}$ $=\frac{1}{T}\sum_{l=1}^{T}\psi^{\text{no}}(t_{l,j}).$
@@
Next, the probability that $x_{i}$ belongs to each class is calculated by computing the cosine similarity between the visual and textual representations and applying softmax.
$p_{ij}=\frac{\exp{(f_{i}\cdot g_{j}/\tau)}}{\sum_{k=1}^{K}\exp{(f_{i}\cdot g_{k}/\tau)}},$ (2)
@@
where $\tau$ is a temperature parameter that adjusts model confidence. The probability that $x_{i}$ does not belong to each class $p_{ij}$ is defined as follows.
$p_{ij}^{\text{no}}=\frac{\exp{(f_{i}\cdot g_{j}^{\text{no}}/\tau)}}{\exp{(f_{i}\cdot g_{j}/\tau)}+\exp{(f_{i}\cdot g_{j}^{\text{no}}/\tau)}}.$ (3)
@@
$p_{ij}^{\text{ID}}$ $=p_{ij}(1-p_{ij}^{\text{no}})$ (4)
$p_{i}^{\text{OOD}}$ $=1-\sum_{j}^{K}{p_{ij}(1-p_{ij}^{\text{no}})}$